OpenStudy (anonymous):

one third the sum of 13 and a certain number is the same as one more than twice the number. find the number

OpenStudy (anonymous):

i got 2

OpenStudy (anonymous):

how

OpenStudy (mathmath333):

(13+x)/3=1+2x

OpenStudy (anonymous):

whats 1/3 times 13

OpenStudy (anonymous):

(13+x)/3=1+2x

OpenStudy (anonymous):

how do u solve this?

OpenStudy (mathmath333):

13+x=3(1+2x)

OpenStudy (mathmath333):

13+x=3+6x

OpenStudy (mathmath333):

13-3=6x-x 10=5x 10/5=x 2=x

OpenStudy (anonymous):

thanks